OpenStudy (anonymous):

solve: 7e^(2x)+2.5=20

OpenStudy (anonymous):

0.4581

OpenStudy (anonymous):

how did you do that?!

OpenStudy (jdoe0001):

can you do something with the 2.5?

OpenStudy (anonymous):

can't you subtract it from 20?

OpenStudy (jdoe0001):

can't? why not?

OpenStudy (anonymous):

ill take that as a yes?

OpenStudy (anonymous):

yes u can and then divide by 7 and then take ln from both sides

OpenStudy (anonymous):

but after that I ge stuck

OpenStudy (jdoe0001):

so can we say that \(\bf 7e^{2x}+2.5=20 \implies 7e^{2x}=17.5\ \ ?\)

OpenStudy (anonymous):

yes, then you divide both sides by 7

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so can we say that \(\bf 7e^{2x}+2.5=20 \implies 7e^{2x}=17.5 \implies 7e^{2x}= \cfrac{17.5}{7}\) ?

OpenStudy (jdoe0001):

woops, darn Imeant \(\bf 7e^{2x}+2.5=20 \implies 7e^{2x}=17.5 \implies e^{2x}= \cfrac{17.5}{7}\) ?

OpenStudy (anonymous):

\[e^(2x) = 2.5\]

OpenStudy (anonymous):

and then I get stuck again

OpenStudy (jdoe0001):

well, ok so \(\bf 7e^{2x}= 2.5\) now let us use the log cancellation rule of => \(\bf log_ab^x= x\) keeping in mind that \(\bf log_ee^x = x \) and that \(\bf log_e = ln\)

OpenStudy (jdoe0001):

well, the rule is really \(\bf \Large log_aa^x= x\) the log base and the "result value" must be equal

OpenStudy (anonymous):

so do you have to plug 2.5 into that somehow?

OpenStudy (jdoe0001):

hmm, actually .. they don't have to be... :( anyhow, that's the cancellation rule

OpenStudy (jdoe0001):

well, we could have used \(\bf log_{10}\) effectively just the same anyhow, I just happen to use ln in this case well what you do is take log to both sides, either \(\bf log_{10}\) or \(\bf ln\) will do :)

OpenStudy (anonymous):

so that would be \[\log_{10} e^(2x) = \log_{10}2.5\]

OpenStudy (jdoe0001):

\(\bf 7e^{2x}+2.5=20 \implies 7e^{2x}=17.5 \implies e^{2x}= 2.5\\ e^{2x}= 2.5 \implies log_{10}(e^{2x}) = log_{10}(2.5)\\ \color{blue}{log_ab^x= x \implies log_{10}(e^{2x}) = 2x}\\ log_{10}(e^{2x}) = log_{10}(2.5) \implies 2x = log_{10}(2.5)\)

OpenStudy (jdoe0001):

hmmm, lemme recheck myself..

OpenStudy (jdoe0001):

hmmm, no I think it has to be the same base as the value, lemme rewrite that using ln :(

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf 7e^{2x}+2.5=20 \implies 7e^{2x}=17.5 \implies e^{2x}= 2.5\\ e^{2x}= 2.5 \implies log_{e}(e^{2x}) = log_{e}(2.5)\\ \color{blue}{log_ab^x= x \implies log_{e}(e^{2x}) = 2x}\\ log_{e}(e^{2x}) = log_{e}(2.5) \implies 2x = log_{e}(2.5)\)

OpenStudy (jdoe0001):

\(\bf \color{blue}{log_aa^x= x \implies log_{e}(e^{2x}) = 2x}\)

OpenStudy (jdoe0001):

so there, 2x = ln(2.5) => x = ln(2.5)/2

OpenStudy (anonymous):

that makes a lot more sense :)

OpenStudy (jdoe0001):

yw